Circular No. 7797

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002D, 2002E, 2002F, and 2002G B. Swift and W. D. Li, University of California at Berkeley, report

observations by LOTOSS (cf. IAUC 7514) of four new apparent supernovae in unfiltered Katzman Automatic Imaging Telescope (KAIT) images:

SN  2002 UT  α_{2000}  δ_{2000}  R  Offset
2002D Jan. 16.2 1^{h}32^{m}57.^{s}80 −16^{o}32′07′′.3 16.8 12′′.6 E, 1′′.6 N
2002E Jan. 16.5 12 08 52.90 − 9 02 13.2 16.8 2′′.2 W, 1′′.7 S
2002F Jan. 17.2 3 52 58.52 +35 35 00.8 18.1 48′′.4 W, 20′′.9 S
2002G Jan. 18.5 13 07 55.26 +34 05 06.5 17.7 5′′.9 E, 7′′.0 S

Additional magnitude estimates from KAIT images: SN 2002D in NGC 594, 2001 Oct. 10.3 UT, [19.5; 2002 Jan. 17.2, 16.7. SN 2002E in NGC 4129, 2001 June 10.3, [19.0; 2002 Jan. 17.5, 16.8. SN 2002F in UGC 2885, 2001 Nov. 9.2, [19.0; 2002 Jan. 18.2, 17.8. SN 2002G, 2001 June 10.3, [19.0; 2002 Jan. 19.4, 17.3.

SUNGRAZING COMETS Further to IAUC 7764, several additional comets have been found on

SOHO website images by S. Hönig (C/2001 X4, X9), M. Meyer (C/2001 X5, X6), M. Boschat (C/2001 X7, Y5), A. Mimeev (C/2001 X8), and X. Zhou (C/2001 Y2, Y3, Y4). Astrometry (measurements by D. Hammer, reduction by B. G. Marsden) and orbital elements appear on the MPECs cited below. All but C/2001 X8 (q ∼ 0.037 AU, i ∼ 72^{o}) appear to be Kreutz sungrazers; all the comets were visible in the C2 coronograph (C/2001 Y5, which was fairly bright with a long tail, was also visible in C3).

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 X4 Dec. 2.396  16^{h}36.^{m}7  −23^{o}47′  2002-A49
C/2001 X5  4.455  16 46.3  −24 08  2002-A49
C/2001 X6  6.246  16 54.1  −24 18  2002-A49
C/2001 X7  7.979  17 04.4  −24 34  2002-A49
C/2001 X8  12.671  17 20.8  −22 17  2002-B01
C/2001 X9  12.814  17 25.0  −25 00  2002-B01
C/2001 Y2  16.463  17 42.6  −25 09  2002-B01
C/2001 Y3  16.521  17 44.0  −24 39  2002-B01
C/2001 Y4  18.246  17 50.2  −25 16  2002-B04
C/2001 Y5  21.529  18 08.4  −28 20  2002-B04
2002 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green